OpenStudy (anonymous):

solve: 2log7-log(x+3)=2

OpenStudy (anonymous):

start by \[\log(7^2)-\log(x+3)=2\] then \[\log(\frac{49}{x+3})=2\]\ and finally write in equivalent exponential form as \[\frac{49}{x+3}=10^2=100\]

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

yw